OpenStudy (anonymous):

Establish the identity 5cos(theta)^2+3sin(theta)^2=3+2cos^2(theta)

OpenStudy (mathmath333):

5cos(theta)^2+3sin(theta)^2 =2cos(theta)^2+(3cos^2(theta)+3sin(theta)^2) =2cos(theta)^2+3(cos^2(theta)+sin(theta)^2) =2cos(theta)^2+3*1

OpenStudy (anonymous):

thanks